OpenStudy (anonymous):

Find the volume of a regular tetrahedron whose altitude measures 10 cm?

OpenStudy (anonymous):

V = sqrt(2) * a^3 / 12 V = sqrt(2) * 1000 / 12 V = sqrt(2) * 500 / 6 V = 250*sqrt(2)/3

jhonyy9 (jhonyy9):

so i think you need to know what mean regular tetrahedron ,yes ?